A firm decides to expand its operations and use more square footage in their main office. Currently, they rent out 3000 square feet of space at a rate of $3,441.00 per month. The new expansion will use this space. The firm has a cost of capital of 9.00% APR. If the tax rate is 38.00% facing the firm, what is the opportunity cost of using this space? (consider this as a perpetuity and express your answer as a positive number)
Answer #1

Given:

Rent per month = $3441

APR = 9%

tax rate = 38%

rent for a year = 3441*12 = $41292

Present value of all the perpetual cashflows

= 41292 / cost of capital

= 41292 / 0.09

=$458800

Tax on these cash flows = 38% * 458800 = $174344

Cash flows after tax = 45880 - 174344 = $284456

A cash flow of $284456 is the opportunity cost of using the space.